OpenStudy (anonymous):

find a general form equation for the line through the pair of points (-1,-5)and(-4,-2)

OpenStudy (anonymous):

One minute. :)

OpenStudy (anonymous):

x + y = -6

OpenStudy (anonymous):

That's your answer. c:

OpenStudy (anonymous):

general form is \[\large ax+by+c=0 \]

OpenStudy (anonymous):

If that is general form, then the equation is: x + y + 6 = 0

OpenStudy (anonymous):

please also show how u got the the answer.... i don't think the asker gained anything from just getting the answer....

OpenStudy (anonymous):

where is the six coming from?

OpenStudy (anonymous):

Using the slope formula, y - y1 / x - x1, I found the slope. -5 - (-2) / (-1 - (-4) -5 + 2 / -1 + 4 -3 / 3 -1 That is the slope. So you can graph the points given, and find more points on the line with the slope. Doing that, I saw that the line intersected the x-axis at (-6, 0). That is where the -6 came from. The standard form of the equation is: x + y = -6 So the general form is: x + y + 6 = 0 Do you understand? :)

OpenStudy (anonymous):

yeah i get it thanks to both of you

OpenStudy (anonymous):

You are welcome. :)